Hello Tutors!I have a problem that I'm stuck with here.
1) The equations of three planes are:
3x + 2y + 7z = 25
2x + y + 4z = 14
5x + 3y + 11z = 39
Solve the above system of equations. State whether this system is consistent or inconsistent and how many solutions are there?

Thank you for your help!!

Answer by stanbon(75887)

The equations of three planes are:
3x + 2y + 7z = 25
2x + y + 4z = 14
5x + 3y + 11z = 39
--------------
The determinant of the coefficient matrix is zero,
so the system is inconsistent. There are no solutions.
=================
Cheers
Stan H.